[figure: 9.7 cm]

CsDy(MoO_{4})_{2} AND KEr(MoO_{4})_{2} - LOW-DIMENSIONAL ISING

MAGNETS

M. Kačmár^{a}, A. Orendáčová^{a}, E. Čižmár^{a}, D. Horváth^{a}, M.Orendáč^{a},

V. Bondarenko^{b}, A.G. Anders^{b}, A. Feher^{a}, M. Jaščur^{a}

aFaculty of Science, P.J. Šafárik University, Park Angelinum 9, 04154 Košice, Slovakia

bInstitute of Low Temp. Physics and Engineering, Lenin Avenue 47, 310164 Kharkov, Ukraine

KEr(MoO_{4})_{2} (I) and CsDy(MoO_{4})_{2} (II) represent low-dimensional magnetic systems with dominant dipolar interactions with ordering temperatures T_{c} ≈ 1 K. Previous studies in zero magnetic field, B, revealed that at helium temperatures (I) resembles a behaviour of S=3/2 rectangular Blume-Capel model with intrachain interaction J_{1}/k_{B}=0.85 K, interchain interaction J_{2}/k_{B} ≈ 0.2 K and single-ion anisotropy 2.D/k_{B}=21 K, while in (II) - S=1/2 Ising system, more intensive strength of interchain correlations was observed, J_{2}/J_{1} ≈ 1, J_{1}/k_{B}=0.6 K. In the present work the behaviour of cluster excitations in B parallel to the easy axis was experimentally studied by the measurements of specific heat, C, vs. B dependence studied at constant temperatures. Corresponding Monte Carlo (MC) simulations of C revealed that despite of the fact that exchange-coupling models proved useful for a sufficient description of the dipolar systems in B=0, the analysis of current experimental data requires a partial renormalization of previously obtained J_{1}, J_{2} parameters. The influence of crystal field effects and long-range character of spin correlations is discussed. Furthermore, MC calculations revealed that the formation of cluster excitations in paramagnetic regime, studied in the chain direction, resembles the behaviour of S=1/2 Ising chain.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : A. Orendáčová

Address for correspondence : Faculty of Science P.J. Šafárik University Park Angelinum 9 04154 Košice, Slovakia

Email address : aorendac@kosice.upjs.sk